# Geometric Representation Condition Improves Equivariant Molecule Generation

### v1

@@ E.2.1 Parametric Conditional Score Matching Error
This is a direct extension of Lemma C.1 from Yuan et al. (2023). Note that we consider the special case where the low-dimensional subspace is the original space (i.e., $A=I_{d_{r}}$ and $d=D=d_{r}$ in their paper), and our noised linear assumption between $r$ and $y$ is identical to their pseudo labeling setting (i.e., $\hat{y}=\hat{w}^{\top}r+\xi$ where $\xi\sim\mathcal{N}(0,\nu^{2})$). We only provide the proof sketch here.
@@ F.1 Representation Visualization
Figure 7: T-SNE visualizations of representations unconditionally generated by the representation generator ($\mathcal{T}=1.0$) vs. those produced by the pre-trained encoder on the QM9 and DRUG datasets. The Silhouette Score is scaled by $\mathbf{10^{2}}$ for clarity.

### v2

@@ E.2.1 Parametric Conditional Score Matching Error
This is a direct extension of Lemma C.1 from (Yuan et al., 2023). Note that we consider the special case where the low-dimensional subspace is the original space (i.e., $A=I_{d_{r}}$ and $d=D=d_{r}$ in their paper), and our noised linear assumption between $r$ and $y$ is identical to their pseudo labeling setting (i.e., $\hat{y}=\hat{w}^{\top}r+\xi$ where $\xi\sim\mathcal{N}(0,\nu^{2})$). We only provide the proof sketch here.
@@ F.1 Representation Visualization
Figure 7: t-SNE visualizations of representations unconditionally generated by the representation generator ($\mathcal{T}=1.0$) vs. those produced by the pre-trained encoder on the QM9 and GEOM-DRUG datasets. The Silhouette Score is scaled by $\mathbf{10^{2}}$ for clarity.